payton:

Bob fenced in his backyard. The perimeter of the yard is 22 feet, and the length of his yard is 5 feet. Use the perimeter formula to find the width of the rectangular yard in inches: P = 2L + 2W. 6 in. 12 in. 72 in. 144 in.

MrMudd183:

i think it's 6. all i did was mutliply 6 by 2 and added that by 10 to get 22, which is the perimeter of the yard

payton:

okay... see i knew you can help

payton:

thats why your my favorite person here

MrMudd183:

thank you UwU

payton:

^_^

MrMudd183:

^w^

SandMandGirl:

First convert 22 feet and 5 feet into inches the put the two values with the formula

SandMandGirl:

22 feet = 2 * 5 + 2 * x 22 = 2 ( 5 + x ) 11 = 5 + x 6 = x 6 feet is 72 inches So 72